You receive a $16,000 5-year constant amortization loan (CAL). The loan's annual interest rate is 10%. What is the total payment in year 4, rounded to the nearest dollar?
Answer #1

Total payment(Principal + Interest) in year 4 is $ 4,221. Annual payment is same for all years.

Annual payment = Loan amount / Present value of annuity of 1
= $       16,000 / 3.790787
= $         4,221
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.10)^-5)/0.10 i = 10%
= 3.79078677 n = 5